OpenStudy (anonymous):

Landon opened a savings account with a deposit of $3,786.45. Nine years later, the balance in the account is $5,135.29. If the interest rate is 3.4%, how often does the interest compound?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (wolf1728):

For a compound interest calculation the formula is total = princ*(1+Rate)^years We know the rate is 3.4 % but we need to know the compounded rate Solving this for rate log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = (log (5,135.29) -log(3,786.45) / 9 log(1 + rate) = (3.7105649741 - 3.5782322263) / 9 log(1 + rate) = 0.1323327477 /9 log(1 + rate) = 0.014703639 1 + rate = 1.0344360295 rate = .0344360295 or 3.44360295 Now we have to determine what compounding rate is needed to change 3.4% to 3.44360295%

OpenStudy (anonymous):

which are annually, daily, quarterly and monthly right?

OpenStudy (wolf1728):

Compounded rate = (1 + [rate/n])^n where n is number of compounding periods per year Compounded rate = (1 + .034/n)^n Solving by trial and error Semi-Annual Compounded rate = (1 + .034)^n Compounded rate = (1.017^2) = 1.034289 rate = 3.4289 % which is lower than 3.44360295% so we need to compound more frequently Quarterly Compounded Rate Compounded rate = (1 + .034/4)^n Compounded rate = (1+ 0.0085) ^ n = 1.034436 rate = 3.4436% So it must be compounded quarterly Yes it's just that simple LOL

OpenStudy (anonymous):

So the Answer is Quarterly?

OpenStudy (wolf1728):

And to answer your question compounding can be done semi-annually, quarterly, monthly, daily, or continuously.

OpenStudy (wolf1728):

Yes the answer is quarterly. Of course if you need to show the math to back up your answer, just copy and paste my 2 postings.

OpenStudy (anonymous):

Thank you for the help!!

OpenStudy (wolf1728):

u r welcome Cjones